OpenStudy (anonymous):

Solve the absolute value equation |3n + 6| = 12. One solution is n = 2. What is the other solution?

myininaya (myininaya):

\[3n+6=-12\] solve this to see

OpenStudy (hoblos):

|3n + 6| = 12 then 3n + 6 =12 OR 3n+6 =-12 n= 2 n=-6

OpenStudy (anonymous):

(3n+6)^2=12^2 => 3n+6=+-12

OpenStudy (anonymous):

On a completely unrelated note, I wonder if FFM would give me a medal for an off-topic response such as this?

OpenStudy (anonymous):

what is FFM?

OpenStudy (anonymous):

FoolForMath. He was viewing the question ;)